# Depth Anywhere: Enhancing 360 Monocular Depth Estimation via Perspective Distillation and Unlabeled Data Augmentation

### v1

@@ 1 Introduction
Our approach consists of two key stages: offline mask generation and online joint training. During the offline stage, we employ a combination of detection and segmentation models to generate masks for invalid regions such as sky and watermarks in unlabeled data. Subsequently, in the online stage, we adopt a semi-supervised learning strategy, loading half of the batch with labeled data and the other half with pseudo-labeled data. Through joint training with both labeled and pseudo-labeled data, our method achieves robust depth estimation performance on 360-degree imagery.
@@ 2 Related Work
  • Directly Apply: Some methods directly apply monocular depth estimation techniques to 360-degree imagery. OmniDepth [55] leverages spherical geometry and incorporates SphConv [31] to improve depth prediction with distortion.  [56, 37] use spherical coordinates to overcome distortion with extra information. SliceNet [21] and ACDNet [54] propose advanced network architectures tailored for omnidirectional images. EGFormer [50] introduces a transformer-based model that captures global context efficiently, while  [33, 32] focuses on integrating geometric priors into the learning process.
  • Cube: Other approaches use cube map projections to mitigate distortion effects. 360-SelfNet [34] is the first work to self-supervised 360 depth estimation leveraging cube-padding [7]. BiFuse [35] and its improved version BiFuse++ [36] are two-branch architectures that utilize cube maps and equirectangular projections. UniFuse [11] combines equirectangular and cube map projections and simplifies the architecture.
  • Tangent Image: Tangent image projections are also popular. 360MonoDepth [26] and OmniFusion [16] convert equirectangular images into a series of tangent images, which are then processed using conventional depth estimation networks. PanoFormer [29] employs a transformer-based architecture to handle tangent images, while SphereNet [9] and HRDFuse [1] enhance depth prediction by collaboratively learning from multiple projections.
@@ 3 Methods
Figure 2: Training Pipeline. Our proposed training pipeline involves joint training on both labeled 360 data with ground truth and unlabeled 360 data. (a) For labeled data, we train our 360 depth model with the loss between depth prediction and ground truth. (b) For unlabeled data, we propose to distill knowledge from a pre-trained perspective-view monocular depth estimator. In this paper, we use Depth Anything [45] to generate pseudo ground truth for training. However, more advanced techniques could be applied. These perspective-view monocular depth estimators fail to produce reasonable equirectangular depth as there exists a domain gap. Therefore, we distill knowledge by inferring six perspective cube faces and passing them through perspective-view monocular depth estimators. To ensure stable and effective training, we propose generating a valid pixel mask with Segment Anything [13] while calculating loss. (c) Furthermore, we augment random rotation on RGB before passing it into Depth Anything, as well as on predictions from the 360 depth model.
In this work, we propose a novel training approach for 360-degree monocular depth estimation models. Our method leverages a perspective depth estimation model as a teacher and generates pseudo labels for unlabeled 360-degree images using a 6-face cube projection. Figure 2 illustrates our training pipeline, incorporating the use of Segment Anything to mask out sky and watermark regions in unlabeled data during the offline stage. Subsequently, we conduct joint training using both labeled and unlabeled data, allocating half of the batch to each. The unlabeled data is supervised using pseudo labels generated by Depth Anything, a state-of-the-art perspective monocular depth foundation model. With the benefit of our teacher model, the 360-degree depth model demonstrates an observable improvement on zero-shot dataset 1.
@@ 3.1 Unleashing the Power of Unlabel 360 data
To tackle the challenges posed by limited data and labeling difficulties in 360-degree datasets, we leverage a large amount of unlabeled data alongside state-of-the-art perspective depth foundation models. Due to significant differences in camera projection and distortion, directly applying perspective models to 360-degree data often yields inferior results. Previous works have explored various methods of projection for converting equirectangular to perspective depth as stated in Sec 2. Among these, cube projection and tangent projection are the most common techniques. We selected cube projection to ensure a larger field of view for each patch, enabling better observation of relative distances between pixels or objects during the inference of the perspective foundation model and enhancing knowledge distillation.
In our approach, we apply projection to unlabeled 360-degree data and then run Depth Anything on these projected patches of perspective images to generate pseudo-labels. We explore two directions for pseudo-label supervision: projecting the patch to equirectangular and computing in the 360-degree domain or projecting the 360-degree depth output from the 360 model to patches and computing in the perspective domain. Since training is conducted in an up-to-scale relative depth manner, stitching the patch of perspective images back to equirectangular with an aligned scale is an additional research aspect. We opt to compute the loss in the perspective domain, facilitating faster and easier training without the need for additional alignment optimization.
@@ 3.2 Random Rotation Processing
Directly applying Depth Anything on cube-projected unlabeled data does not yield improvements due to ignorance of cross cube-face relation, leading to cube artifacts (Figure 4). This issue arises from the separate estimation of perspective cube faces, where monocular depth is estimated based on semantic information, potentially lacking a comprehensive understanding of the entire scene. To address this, we propose a random rotation preprocessing step in front of the perspective foundation model.
@@
Figure 4: Cube Artifact. As shown in the center row of the figure, an undesired cube artifact appears when we apply joint training with pseudo ground truth from Depth Anything [45] directly. This issue arises from independent relative distances within each cube face caused by a static point of view. Ignoring cross-cube relationships results in poor knowledge distillation. To address this, as shown in Figure 2 (c), we randomly rotate the RGB image before inputting it into Depth Anything. This enables better distillation of depth information from varying perspectives within the equirectangular image.
@@ 4.1 Baselines
Recent state-of-the-art methods [1, 50, 35, 36, 11, 21, 29] have emerged. We chose UniFuse and BiFuse++ as our baseline models for experiments, as many of the aforementioned methods did not release pre-trained models or provide training code and implementation details. It’s worth noting that PanoFormer [29] is not included due to incorrect evaluation code and results, and EGFormer [29] is not included since its experiments are mainly conducted on other datasets and benchmarks. Both selected models are re-implemented with affine-invariant loss on disparity for a fair comparison and to demonstrate improvement. We conduct experiments on the Matterport3D benchmark to demonstrate improvement within the same dataset/domain and apply zero-shot evaluation on the Stanford2D3D test set to demonstrate the generalization ability of our proposed training technique.
@@ 4.2 Benchmarks evaluation
BiFuse BerHu M M - 0.8452 0.9319 0.9632
UniFise BerHu M M 0.1063 0.8897 0.9623 0.9831
SliceNet BerHu M M - 0.8716 0.9483 0.9716
BiFuse++ BerHu M M - 0.879 0.9517 0.9772
HRDFuse BerHu M M 0.0967 0.9162 0.9669 0.9844
UniFuse Affine-Inv M M 0.10187 0.89311 0.97044 0.98902
UniFuse Affine-Inv M, ST-all (p) M 0.08938 0.91143 0.97512 0.99078
BiFuse++ Affine-Inv M M 0.0938 0.9142 0.97363 0.9894
BiFuse++ Affine-Inv M, ST-all (p) M 0.085 0.917 0.976 0.991
@@ 4.3 Zero-shot evaluation
Depth Anything [45] and Marigold [12] are state-of-the-art zero-shot depth models trained with perspective depths. As shown in Table 3, due to the domain gap and different camera projections, foundation models trained with perspective depth cannot be directly applied to 360-degree images. We demonstrated the zero-shot improvement on UniFuse [11] and BiFuse++ [36] with models trained on the entire Matterport3D [5] dataset with ground truth and the entire Structured3D [51] or SpatialAudioGen [20] dataset with pseudo ground truth generated using Depth Anything [45].
@@
Table 3: Zero-shot Evaluation on Stanford2D3D. We perform zero-shot evaluations with models trained on other datasets. Following BiFuse and UniFuse’s settings, we train the 360 models on the entire Matterport3D dataset and then test on Stanford3D’s test set.
@@
BiFuse BerHu M-all SF 0.1195 0.8616 - -
UniFuse BerHu M-all SF 0.0944 0.9131 - -
BiFuse++ BerHu M-all SF 0.107 0.914 0.975 0.989
Depth Anything Affine-Inv Pers. SF 0.248 0.635 0.899 0.97
Marigold Affine-Inv Pers. SF 0.195 0.692 0.942 0.982
UniFuse Affine-Inv M-all SF 0.09 0.914 0.976 0.99
UniFuse Affine-Inv M-all, ST-all (p) SF 0.086 0.924 0.977 0.99
UniFuse Affine-Inv M-all, SP-all (p) SF 0.09 0.92 0.978 0.99
BiFuse++ Affine-Inv M-all SF 0.09 0.921 0.976 0.99
BiFuse++ Affine-Inv M-all, ST-all (p) SF 0.082 0.931 0.979 0.991
BiFuse++ Affine-Inv M-all, SP-all (p) SF 0.858 0.926 0.979 0.991
Figure 5: Zero-shot qualitative with UnFuse [11] (left 2 columns) and BiFuse++ [36] (right 2 columns) tested on Stanford2D3D.
@@
UniFuse Affine-Inv M-all ST 0.202 0.759 0.932 0.97
UniFuse Affine-Inv M-all, ST-all (p) ST 0.13 0.887 0.953 0.977
UniFuse Affine-Inv M-all, SP-all (p) ST 0.152 0.864 0.946 0.972
@@ 4.4 Qualtative results in the wild
Figure 6: Generalization ability in the wild with depth map visualization: We showcase zero-shot qualitative results using a combination of images captured by us and randomly sourced from the internet to assess the model’s generalization ability. For privacy reasons, we have obscured the camera-man in the images.
Figure 7: Generalization ability in the wild with point cloud visualization: We showcase zero-shot qualitative results in point cloud using a combination of images captured by us and randomly sourced from the internet to assess the model’s generalization ability.
@@ 5 Conclusion
Limitations
Our work faces limitations due to its heavy reliance on the quality of unlabeled data and pseudo labels from perspective foundation models. The results are significantly impacted by data quality (Sec. 3.1). Without data cleaning, the training process resulted in NaN values.
@@ A.2 More qualitative
We demonstrate additional zero-shot qualitative results in Figure 8. In-domain results on the Matterport3D test sets are showcased in Figure 9 and Figure 10.
@@ A.3 Dataset statistic
Table 5: 360 monocular depth estimation lacks a large amount of training data. This table lists datasets used in 360-degree monocular depth estimation alongside perspective depth datasets from the Depth Anything methodology. The volume of training data for 360-degree imagery (right column) is significantly smaller than that for perspective imagery (left column), by about 200 times. This highlights the need for using perspective distillation techniques to enhance the limited data available for 360-degree depth estimation. Ground truth (GT) labels are noted where applicable, showing the available resources for training in these domains.

### v2

@@ 1 Introduction
Our approach consists of two key stages: offline mask generation and online joint training. During the offline stage, we employ a combination of detection and segmentation models to generate masks for invalid regions, such as sky and watermarks in unlabeled data. Subsequently, in the online stage, we adopt a semi-supervised learning strategy, loading half of the batch with labeled data and the other half with pseudo-labeled data. Through joint training with both labeled and pseudo-labeled data, our method achieves robust depth estimation performance on 360-degree imagery.
@@ 2 Related Work
  • Directly Apply: Some methods directly apply monocular depth estimation techniques to 360-degree imagery. OmniDepth [69] leverages spherical geometry and incorporates SphConv [42] to improve depth prediction with distortion.  [70, 49] use spherical coordinates to overcome distortion with extra information. [12, 17] leverage other ground truth supervisions to assist on depth estimation. SliceNet [31] and ACDNet [68] propose advanced network architectures tailored for omnidirectional images. EGFormer [64] and HiMODE [18] introduce a transformer-based model that captures global context efficiently, while  [45, 44] focuses on integrating geometric priors into the learning process.  [13] proposed to generate large-scale datasets with SfM and MVS, then apply to test-time training.
  • Cube: Other approaches use cube map projections to mitigate distortion effects. 360-SelfNet [46] is the first work to self-supervised 360 depth estimation leveraging cube-padding [9]. BiFuse [47] and its improved version BiFuse++ [48] are two-branch architectures that utilize cube maps and equirectangular projections. UniFuse [16] combines equirectangular and cube map projections and simplifies the architecture.  [3] combines two-branch techniques with transformer network.
  • Tangent Image: Tangent image projections are also popular. [37, 23, 30] convert equirectangular images into a series of tangent images, which are then processed using conventional depth estimation networks. PanoFormer [40] employs a transformer-based architecture to handle tangent images, while SphereNet [11] and HRDFuse [1] enhance depth prediction by collaboratively learning from multiple projections.
360 other works
Beyond depth estimation, 360-degree imagery has been applied to depth completion tasks as follows [26, 8, 32, 57, 58, 15]. Other methods, such as  [25] and  [43] focus on the projection between camera models, while the former projects pinhole camera model images into a large field of view, whereas the latter transforms convolution kernels.
@@ 3 Methods
Figure 2: Training Pipeline. Our proposed training pipeline involves joint training on both labeled 360 data with ground truth and unlabeled 360 data. (a) For labeled data, we train our 360 depth model with the loss between depth prediction and ground truth. (b) For unlabeled data, we propose to distill knowledge from a pre-trained perspective-view monocular depth estimator. In this paper, we use Depth Anything [59] to generate pseudo ground truth for training. However, more advanced techniques could be applied. These perspective-view monocular depth estimators fail to produce reasonable equirectangular depth as a domain gap exists. Therefore, we distill knowledge by inferring six perspective cube faces and passing them through perspective-view monocular depth estimators. To ensure stable and effective training, we propose generating a valid pixel mask with Segment Anything [20] while calculating loss. (c) Furthermore, we augment random rotation on RGB before passing it into Depth Anything, as well as on predictions from the 360 depth model.
In this work, we propose a novel training approach for 360-degree monocular depth estimation models. Our method leverages a perspective depth estimation model as a teacher and generates pseudo labels for unlabeled 360-degree images using a 6-face cube projection. Figure 2 illustrates our training pipeline, incorporating the use of Segment Anything to mask out sky and watermark regions in unlabeled data during the offline stage. Subsequently, we conduct joint training using both labeled and unlabeled data, allocating half of the batch to each. The joint training avoids limiting performance by teacher model. The unlabeled data is supervised using pseudo labels generated by Depth Anything, a state-of-the-art perspective monocular depth foundation model. With the benefit of our teacher model, the 360-degree depth model demonstrates an observable improvement on the zero-shot dataset, as shown in Figure 1.
@@ 3.1 Unleashing the Power of Unlabel 360 data
To tackle the challenges posed by limited data and labeling difficulties in 360-degree datasets, we leverage a large amount of unlabeled data alongside state-of-the-art perspective depth foundation models. Due to significant differences in camera projection and distortion, directly applying perspective models to 360-degree data often yields inferior results. Previous works have explored various methods of projection for converting equirectangular to perspective depth, as stated in Section 2. Among these, cube projection and tangent projection are the most common techniques. We selected cube projection to ensure a larger field of view for each patch, enabling better observation of relative distances between pixels or objects during the inference of the perspective foundation model and enhancing knowledge distillation. The comparison table can be find in the supplementary material.
In our approach, we apply projection to unlabeled 360-degree data and then run Depth Anything on these projected patches of perspective images to generate pseudo-labels. We explore two directions for pseudo-label supervision: projecting the patch to equirectangular and computing in the 360-degree domain or projecting the 360-degree depth output from the 360 model to patches and computing in the perspective domain. Since training is conducted in an up-to-scale relative depth manner, stitching the patch of perspective images back to equirectangular with an aligned scale will lead to failure in training Figure 4, which is an additional research topic that is worth investigation. We opt to compute the loss in the perspective domain, facilitating faster and easier training without the need for additional alignment optimization.
Figure 4: Qualitative visualization of a model trained directly on pseudo equirectangular data without scale alignment. We propose calculating the loss with pseudo ground truth on cube faces due to scale misalignment between the six faces during the cube-to-equirectangular projection. We showcase the results of a model trained on pseudo equirectangular data without scale alignment as a simple baseline to demonstrate the importance of calculating loss separately on each of the six faces. The images are presented from top to bottom as follows: (a) RGB images. (b) Pseudo cube ground truth projected directly to equirectangular. (c) Prediction trained with row 2. (d) Pseudo cube ground truth with rotation projected directly to equirectangular. (e) Prediction trained with row 4. (f) Our model’s predictions are trained on cube faces separately with rotation.
@@ 3.2 Random Rotation Processing
Directly applying Depth Anything on cube-projected unlabeled data does not yield improvements due to ignorance of cross-cube-face relation, leading to cube artifacts (Figure 5). This issue arises from the separate estimation of perspective cube faces, where monocular depth is estimated based on semantic information, lacking a comprehensive understanding of the entire scene. To address this, we propose a random rotation preprocessing step in front of the perspective foundation model.
@@
Figure 5: Cube Artifact. Shown in the center row of the figure, an undesired cube artifact appears when we apply joint training with pseudo ground truth from Depth Anything [59] directly. This issue arises from independent relative distances within each cube face caused by a static point of view. Ignoring cross-cube relationships results in poor knowledge distillation. To address this, as shown in Figure 2(c), we randomly rotate the RGB image before inputting it into Depth Anything. This enables better distillation of depth information from varying perspectives within the equirectangular image.
@@ 4.1 Baselines
Recent state-of-the-art methods [1, 64, 47, 48, 16, 31, 40] have emerged. We chose UniFuse and BiFuse++ as our baseline models for experiments, as many of the aforementioned methods did not fully release pre-trained models or provide training code and implementation details. It’s worth noting that PanoFormer [40] is not included due to incorrect evaluation code and results. Both selected models are re-implemented with affine-invariant loss on disparity for a fair comparison and to demonstrate improvement. We conduct experiments on the Matterport3D [6] benchmark to demonstrate performance gains within the same dataset/domain, and we perform zero-shot evaluation on the Stanford2D3D [2] test set to demonstrate the generalization capability of our proposed training technique. To further validate its robustness, we evaluate additional baseline models [45, 64] in zero-shot setting, showcasing the effectiveness of our approach for non-dual-projection models.
@@ 4.2 Benchmarks Evaluation
BiFuse [47] BerHu M M - 0.845 0.932 0.963
UniFuse [16] BerHu M M 0.106 0.890 0.962 0.983
SliceNet [31] BerHu M M - 0.872 0.948 0.972
BiFuse++ [48] BerHu M M - 0.879 0.952 0.977
HRDFuse [1] BerHu M M 0.097 0.916 0.967 0.984
UniFuse [16] Affine-Inv M M 0.102 0.893 0.970 0.989
UniFuse [16] Affine-Inv M, ST-all (p) M 0.089 0.911 0.975 0.991
BiFuse++ [48] Affine-Inv M M 0.094 0.914 0.974 0.989
BiFuse++ [48] Affine-Inv M, ST-all (p) M 0.085 0.917 0.976 0.991
@@ 4.3 Zero-Shot Evaluation
Depth Anything [59] and Marigold [19] are state-of-the-art zero-shot depth models trained with perspective depths. As shown in Table 3, due to the domain gap and different camera projections, foundation models trained with perspective depth cannot be directly applied to 360-degree images. We demonstrated the zero-shot improvement on UniFuse [16], BiFuse++ [48] and non-dual-projection methods [45, 64] with models trained on the entire Matterport3D [6] dataset with ground truth and the entire Structured3D [65] or SpatialAudioGen [29] dataset with pseudo ground truth generated using Depth Anything [59].
@@
Table 3: Zero-shot Evaluation on Stanford2D3D. We perform zero-shot evaluations with models trained on other datasets. Following the original training settings, we train the 360 models [48, 16, 45, 64] on the entire Matterport3D dataset and then test on Stanford3D’s test set.
@@
BiFuse [47] BerHu M-all SF 0.120 0.862 - -
UniFuse [16] BerHu M-all SF 0.094 0.913 - -
BiFuse++ [48] BerHu M-all SF 0.107 0.914 0.975 0.989
Depth Anything [59] Affine-Inv Pers. SF 0.248 0.635 0.899 0.97
Marigold [19] Affine-Inv Pers. SF 0.195 0.692 0.942 0.982
UniFuse [16] Affine-Inv M-all SF 0.090 0.914 0.976 0.990
UniFuse [16] Affine-Inv M-all, ST-all (p) SF 0.086 0.924 0.977 0.990
UniFuse [16] Affine-Inv M-all, SP-all (p) SF 0.090 0.920 0.978 0.990
BiFuse++ [48] Affine-Inv M-all SF 0.090 0.921 0.976 0.990
BiFuse++ [48] Affine-Inv M-all, ST-all (p) SF 0.082 0.931 0.979 0.991
BiFuse++ [48] Affine-Inv M-all, SP-all (p) SF 0.086 0.926 0.979 0.991
HoHoNet [45] Affine-Inv M-all SF 0.095 0.906 0.975 0.991
HoHoNet [45] Affine-Inv M-all, ST-all (p) SF 0.088 0.920 0.979 0.992
EGFormer [64] Affine-Inv M-all SF 0.098 0.906 0.972 0.989
EGFormer [64] Affine-Inv M-all, ST-all (p) SF 0.086 0.923 0.976 0.990
Figure 6: Zero-shot qualitative with UniFuse [16] (left) and BiFuse++ [48] (right) tested on Stanford2D3D.
@@
UniFuse [16] Affine-Inv M-all ST 0.202 0.759 0.932 0.970
UniFuse [16] Affine-Inv M-all, ST-all (p) ST 0.130 0.887 0.953 0.977
UniFuse [16] Affine-Inv M-all, SP-all (p) ST 0.152 0.864 0.946 0.972
@@ 4.4 Qualtative Results in the Wild
Table 5: Metric depth fine-tuning. We fine-tune our model trained with Matterport3D [6] ground truth label and Structured3D [65] pseudo label on relative depth with Stanford2D3D [2]’s training set metric depths with a single epoch.
Method MAE ↓ Abs Rel ↓ RMSE ↓ RMSElog ↓ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
UniFuse [16] 0.208 0.111 0.369 0.072 0.871 0.966 0.988
UniFuse [16] (Ours) 0.206 0.118 0.351 0.049 0.910 0.971 0.987
Figure 7: Generalization ability in the wild with depth map visualization. We showcase zero-shot qualitative results using a combination of images we captured and randomly sourced from the internet to assess the model’s generalization ability. For privacy reasons, we have obscured the cameraman in the images.
Figure 8: Generalization ability in the wild with point cloud visualization. We showcase zero-shot qualitative results in the point cloud using a combination of images we captured and randomly sourced from the internet to assess the model’s generalization ability.
4.5 Fine-Tuned to Metric Depth Estimation
We use our pre-trained model as an initial weight and fine-tune on Stanford2D3D [2] metric depth to demonstrate the effectiveness of our pre-trained relative depth model’s ability to adapt to metric depth with a single epoch in Table 5
@@ 5 Conclusion
Limitations.
Our work faces limitations due to its heavy reliance on the quality of unlabeled data and pseudo labels from perspective foundation models. The results are significantly impacted by data quality (Section 3.1). Without data cleaning, the training process resulted in NaN values. Another limitation is that although with unlabeled data, the scarcity of data still exists compared to other tasks.
@@ A.2 More Qualitative
We demonstrate additional zero-shot qualitative results in Figure 9 and in-the-wild results in Figure 12. In-domain results on the Matterport3D test sets are showcased in Figure 10 and Figure 11.
@@ A.3 Dataset Statistic
A.4 Ground Truth and Pseudo Label Ratio Ablation
Unlike many previous knowledge distillation approaches that use a higher proportion of pseudo labels during model training, we opt for an equal ratio of ground truth to pseudo labels. Through an ablation study exploring the relationship between this data ratio and model performance, we observe a robust improvement starting from $1:1$ in Table 7.
A.5 Perspective Camera Projection Ablation
There are various perspective camera projections for panoramic imagery, with cube and tangent image projections being the most common, both widely used in previous works. In Table 8, we compare these two projections and observe similar improvements when applying our proposed training pipeline, demonstrating the effectiveness of our method. For robust knowledge distillation in relative depth estimation, we select the cube projection due to its wider field of view coverage.
Table 6: 360 monocular depth estimation lacks a large amount of training data. This table lists datasets used in 360-degree monocular depth estimation alongside perspective depth datasets from the Depth Anything methodology. The volume of training data for 360-degree imagery (right column) is significantly smaller than that for perspective imagery (left column) by about 200 times. This highlights the need for using perspective distillation techniques to enhance the limited data available for 360-degree depth estimation. Ground truth (GT) labels are noted where applicable, showing the available resources for training in these domains.
@@ A.5 Perspective Camera Projection Ablation
Table 7: Ratios of GT and Pseudo label during training. We conduct additional experiments with varying ratios, which shows our method is robust across different ratios starting from 1:1.
Ratio train(GT) train(Pseudo) test Abs Rel ↓ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
1:1 M-all ST-all(p) SF 0.086 0.924 0.977 0.990
1:2 M-all ST-all(p) SF 0.087 0.923 0.977 0.990
1:4 M-all ST-all(p) SF 0.085 0.923 0.977 0.990
Table 8: Comparison between Tangent Image and Cube Projection. We compare two of the most commonly used perspective camera projections for panorama images. As shown in the table, both projections yield similar quantitative results. However, we select the cube projection for knowledge distillation due to its broader field of view coverage.
Projection Method train test Abs Rel ↓ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
Cube UniFuse M-all+ST-all(p) SF 0.086 0.924 0.977 0.990
Tangent UniFuse M-all+ST-all(p) SF 0.087 0.923 0.978 0.991
Figure 12: Additional in-the-wild results. We compare our proposed joint-training method (Matterport3D (GT) + SpatialAudioGen (Pseudo)) with a model trained only on the Matterport3D dataset, using data randomly downloaded from the internet. This comparison demonstrates the significant improvement of our method along with its generalization ability and effectiveness on real-world data.